OpenStudy (anonymous):

20g of a radiioactive substance reduces to 2.5g in 2079 day. What is the half life nd decay constant of the substance?

OpenStudy (anonymous):

srinidhi

OpenStudy (anonymous):

srinidhi plz explain.

OpenStudy (anonymous):

srinidhi i have done.....

OpenStudy (anonymous):

ok wt

OpenStudy (anonymous):

T*k=2.303logA/A-XFROM THIS EQUATION U GET K AFTER THAT U HAVE TO FIND THE HALF LIFE JUST FETCH THE VALUE OF K IN THE EQUATION OF HALF LIFT K=0.693/TIME 4 HALF LIFE

OpenStudy (anonymous):

time= half life x {log(intial amt/final amount) / log 2}

OpenStudy (anonymous):

SRINDHI TUMKO SMJ ME AA GYA

OpenStudy (anonymous):

Half life= 2079 / {(log20/2.5)/log2)} =2079/(log8/log2)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

kya main galat hun kya...agar haan to tell

OpenStudy (anonymous):

kaha

OpenStudy (anonymous):

phele apko K ki value nikalni hogi .......k stands 4 rate contant ok.........after that usi k se tum half life nikal skte ho .......agr nhi smj me aaya to fir puch lena ok......me tumhe equation b bta skta hu

OpenStudy (anonymous):

srinidhi r u there?

OpenStudy (anonymous):

yes, i ws posting another question